THE GEOMETRIC CHANGE OF HF-CO_{2} UPON VIBRATIONAL EXCITATION

\underline{ZHENHONG YU}, TODD A. STONE, CHENG-CHI CHUANG, WALTER DRISDELL, and WILLIAM KLEMPERER, Harvard University, Department of Chemistry and Chemical Biology, Cambridge, Massachusetts, 02138.

HF-CO_{2} is quasilinear at the vibrational ground state^{a} and becomes a semi-rigid bent species at ν_{HF} = 3 with the CO_{2} intermolecular bending frequency of 24.75(10) cm^{−1}. The ab initio potential surface is very flat from θ = 0 to 40^{∘} at ν_{HF} = 0 while at ν_{HF} = 3 the potential energy of linear geometry is about 50 cm^{−1} higher than that at the minimum near θ = 40^{∘}. The observed features at 11174.45, 11168.10, and 11181.82 cm^{−1} have been assigned to parallel transition to the K = 0, 1, and 2 levels of the second overtone HF valence band, respectively, indicating a vibrational redshift of 198.36(5) cm^{−1} compared to the HF monomer. The relatively strong transition intensities of the K subbands are due to the inertial axes switching^{b}. The rotational constants of the (3000000) state are A = 2.96(2) cm^{−1}, (B+C)/2 = 0.0742(10), 0.0717(10), and 0.0696(10) cm^{−1} for the K = 0, 1, and 2 levels. The centrifugal distortion D_{K} = 0.270(5) cm^{−1} is extremely large but in good agreement with the expectation, as a result of very soft CO_{2} intermolecular bending. The spectral linewidths are 9.0(9), 7.2(6), and 4.5(6) GHz for the above levels, showing dramatic dependence of vibrational predissociation lifetime upon K. A perpendicular transition of the HF bending combination band at 11538.92 cm^{−1} provides a bending frequency of 362.77(15) cm^{−1}, while (B+C)/2 = 0.0668(10) cm^{−1} and Γ = 4.2(6) GHz.

aD. J. Nesbitt and C. M. Lovejoy, J. Chem. Phys. 93, 7716(1990).

bJ. T. Hougen and J. K. Watson, J. Mol. Spectrosc. 43, 298(1965).